# The Catechol Benchmark: Time-series Solvent Selection Data for Few-shot Machine Learning

### v1

@@ 3.3 Gaussian process extensions
Kernel design must be performed carefully when creating GP models. A key issue with using the standard RBF kernel is that, for unseen solvents with featurizations very dissimilar to the solvents in the train set, the GP will revert to the uninformative prior. We therefore propose decomposing the kernel into solvent and non-solvent components in an additive manner, similarly to Ru et al. [58].
Multitask kernels are able to learn correlations between outputs [59]. For example, the yields of the two products tend to be positively correlated with each other, and negatively correlated with the remaining starting material.
Nonstationary approaches allow modeling of search spaces that have changing lengthscale. For example, the rate of reaction is fastest in the first few minutes, and the solvent mixing may be nonlinear in the feature space as noted in Section 3.1. We therefore learn a warping of these two inputs, inspired by Snoek et al. [60] and Balandat et al. [61].
The results of these extensions are presented in Table 3. These show promising directions to improve regression on single solvents, but struggle to beat the simpler GP when introducing solvent ramps. We encourage the machine learning community to investigate further extensions, such as considering the heteroskedastic nature of the observation noise [62], or non-stationary kernels [63, 64].
@@ 3.4 Transfer learning
A key challenge in this dataset is the relatively low amount of data: where many modern ML approaches require large volumes of data, we only have 1227 observations of reaction conditions. To address this, we extend the training data by including results from the Ethyl dataset [28], which has a further 283 experiments. For this regression problem, we only predict the total product yield, since the Ethyl dataset only has one observed product.
@@
Catechol Catechol + Claisen
@@ 3.5 Active learning and Bayesian optimization
One key application for machine learning in chemistry is to optimally design experimental conditions, with recent interest in transient flow applications [65, 66, 67]. In this section, we showcase how the dataset can be used to benchmark algorithms for design of experiments. For simplicity, we focus on the independent GP model with descriptors from Spange et al. [40]. We first explore active learning ideas in transient flow: we split the dataset into ramps and use the entropy and mutual information criteria [68] to select transient ramps sequentially, with the goal of maximizing information across the dataset, which we measure by MSE. Figure 5(a) shows that using the entropy criterion reduces MSE more initially, while using mutual information gives the best long-term performance.
We then benchmark on classical Bayesian optimization algorithms Expected Improvement [69] and Upper Confidence Bound [70]. We design an objective to maximize product yield and the selectivity of Product 2, while minimizing temperature and residence time, exemplifying conflicting objectives in the scale-up process. We allow the algorithms to query single points across the whole dataset, with the goal of identifying the optimal configuration in the fewest queries. Figure 5(b) shows the results. In this case, we observe that the algorithms are very quickly able to identify the optimum, usually after 20 iterations, outperforming a random search by a significant margin.
@@ A.2.1 Gaussian processes
We implement the GP models in this paper in BoTorch v0.13.0 [61]. We use the priors recommended by Hvarfner et al. [71], to ensure good performance across featurizations of different dimensions. We use an RBF kernel, with the lengthscale prior
@@
Decomposed kernel. We take inspiration from Ru et al. [58], and separate our kernel into two parts. Specifically, we consider the input to the model to be the concatenation of the solvent featurization, $f$, and the non-featurized inputs, $x$, which include residence time and temperature. We then use the following kernel,
@@
Input warping. In Section 3.3, we describe how the underlying chemistry is nonstationary. To attempt to address this, we take inspiration from Snoek et al. [60] and Balandat et al. [61], learning a bijective map $\phi:[0,1]\to[0,1]$ that can capture the nonlinear effect of mixing solvents. This map has hyperparameters that can be learned,
@@ A.3 Additional results
We showcase additional results for Neural Processes [54] and graph Gaussian processes [50, 51] in table 5.
Table 5: Regression performance on the single solvent dataset. Mean squared error (MSE) and negative log predictive density (NLPD) are averaged across all leave-one-out data splits. We include the shortest path kernel (sp) and the encoded shortest path kernel (esp).
@@ B.1 Reactor details
The automated reactor setup used to collect the data is shown in Figure 7. Knauer Azure 4.1S pumps fitted with stainless steel 10 mL pump heads were used as pumps 1 and 2. All tubing used for the entire reactor was made of 316 stainless steel (1.5875 mm OD, 1 mm ID). An Agilent inline jet weaver HPLC mixer (350 $\mu$L volume) was used as an inline mixer to ensure the reactant solution was homogeneous before entering the reactor. An Agilent 6890 GC oven was used to heat the stainless steel coiled reactor (1.5875 mm OD, 1 mm ID, 7.95 mL volume) during the reaction to the desired temperature. A customized cooling system made from an aluminum block and a Peltier assembly was then placed inline to rapidly cool the flow of solution and quench the reaction. A Vici four port-2 position sampling valve followed the Peltier to sample small aliquots (500 nL) into the HPLC for online analysis measurements of the reaction. An IDEX 1000 PSI BPR was then placed before the waste tubing of the reactor to depressurize the reaction solution back to atmospheric pressure. The pumps, oven and Vici valve were automated by code developed in house in Python.
@@ B.1.1 Methods
    The reactant solutions were made up by adding allyl phenyl ether (50 $\mu$L) and the internal standard - ethyl benzene (50 $\mu$L) in to both solvent A and solvent B (250 mL) in separate volumetric flasks.
@@ B.2 Fine-tuning calibration via optimization
[…]
@@
$c_{i}=c_{i^{\prime}}\quad\quad\text{ if }i,i^{\prime}\text{ are in the same experimental run }$
@@
[…]
@@ B.3 Spange descriptor interpolation
The descriptors from Spange et al. [40] were obtained from the supplementary material on the paper. However, there are a few values missing from some rows, including for the solvents we gathered data for. In order to estimate the missing values, we trained a multi-task Gaussian process model on the whole table, under a Taniamoto kernel, which we then used to predict the missing values that are used for all the main methods in the paper.

### v2

@@ 3.3 Gaussian process extensions
Kernel design must be performed carefully when creating GP models. A key issue with using the standard RBF kernel is that, for unseen solvents with featurizations very dissimilar to the solvents in the train set, the GP will revert to the uninformative prior. We therefore propose decomposing the kernel into solvent and non-solvent components in an additive manner, similarly to ru2020bayesian.
Multitask kernels are able to learn correlations between outputs [bonilla2007multi]. For example, the yields of the two products tend to be positively correlated with each other, and negatively correlated with the remaining starting material.
Nonstationary approaches allow modeling of search spaces that have changing lengthscale. For example, the rate of reaction is fastest in the first few minutes, and the solvent mixing may be nonlinear in the feature space as noted in Section 3.1. We therefore learn a warping of these two inputs, inspired by snoek2014input and balandat2020botorch.
The results of these extensions are presented in Table 4. These show promising directions to improve regression on single solvents, but struggle to beat the simpler GP when introducing solvent ramps. We encourage the machine learning community to investigate further extensions, such as considering the heteroskedastic nature of the observation noise [le2005heteroscedastic], or non-stationary kernels [van2017convolutional, boyne2025bark].
@@ 3.4 Transfer learning
A key challenge in this dataset is the relatively low amount of data: where many modern ML approaches require large volumes of data, we only have 1227 observations of reaction conditions. To address this, we extend the training data by including results from the Ethyl dataset [WO2025073762], which has a further 283 experiments. For this regression problem, we only predict the total product yield, since the Ethyl dataset only has one observed product.
@@
Catechol Catechol + Ethyl
@@ 3.5 Active learning and Bayesian optimization
One key application for machine learning in chemistry is to optimally design experimental conditions, with recent interest in transient flow applications [folch2022snake, mutny2023active, folch2024transition]. In this section, we showcase how the dataset can be used to benchmark algorithms for design of experiments. For simplicity, we focus on the independent GP model with descriptors from spange2021reappraisal. We first explore active learning ideas in transient flow: we split the dataset into ramps and use the entropy and mutual information criteria [krause2007nonmyopic] to select transient ramps sequentially, with the goal of maximizing information across the dataset, which we measure by MSE. Figure 5(a) shows that using the entropy criterion reduces MSE more initially, while using mutual information gives the best long-term performance.
We then benchmark on classical Bayesian optimization algorithms Expected Improvement [mockus1998application] and Upper Confidence Bound [srinivas2009gaussian]. We design an objective to maximize product yield and the selectivity of Product 2, while minimizing temperature and residence time, exemplifying conflicting objectives in the scale-up process. We allow the algorithms to query single points across the whole dataset, with the goal of identifying the optimal configuration in the fewest queries. Figure 5(b) shows the results. In this case, we observe that the algorithms are very quickly able to identify the optimum, usually after 20 iterations, outperforming a random search by a significant margin.
Finally we consider a multi-objective optimization benchmark, by considering a three dimensional objective function of trying to optimize yield, selectivity, and a green-score created from Table 2. We consider the solvent greenness by setting a value of 1.0 to every green solvent, -1.0 to every harmful solvent, and 0.0 to the remaining. For mixture solvent data-points we take a weighted average of their green scores. As benchmark metrics we consider three metrics of Pareto coverage: Euclidean generational distance (GD), inverted generational distance (IGD), and the maximum Pareto frontier error (MPFE) [riquelme2015performance]. We present results in Table 6, where we benchmark Thompson Sampling with random scalarazations [paria2020flexible].
@@
Table 6: Results of multi-objective optimization benchmarking over iterations. We compare Euclidean generational distance (GD), inverted generational distance (IGD), and the maximum Pareto frontier error (MPFE). Multi-objective optimization results in much stronger metrics that randomly sampling.
Iteration GD ($\downarrow$) IGD($\downarrow$) MPFE ($\downarrow$)
Random MOBO Random MOBO Random MOBO
1 0.199 0.158 0.397 0.356 0.453 0.387
25 0.107 0.043 0.177 0.146 0.421 0.225
50 0.081 0.023 0.126 0.105 0.388 0.212
75 0.069 0.017 0.106 0.084 0.385 0.200
100 0.054 0.014 0.092 0.072 0.285 0.200
Figure 6: Visualization of empirical 3-dimensional Pareto front for the multi-objective Bayesian optimization benchmark.
@@ A.2.1 Gaussian processes
We implement the GP models in this paper in BoTorch v0.13.0 [balandat2020botorch]. We use the priors recommended by hvarfner2024vanilla, to ensure good performance across featurizations of different dimensions. We use an RBF kernel, with the lengthscale prior
@@
Decomposed kernel. We take inspiration from ru2020bayesian, and separate our kernel into two parts. Specifically, we consider the input to the model to be the concatenation of the solvent featurization, $f$, and the non-featurized inputs, $x$, which include residence time and temperature. We then use the following kernel,
@@
Input warping. In Section 3.3, we describe how the underlying chemistry is nonstationary. To attempt to address this, we take inspiration from snoek2014input and balandat2020botorch, learning a bijective map $\phi:[0,1]\to[0,1]$ that can capture the nonlinear effect of mixing solvents. This map has hyperparameters that can be learned,
@@ A.3 Additional results
We showcase additional results for Neural Processes [garnelo2018conditional] and graph Gaussian processes [borgwardt2005shortest, xie2025bogrape] in table 7.
Table 7: Regression performance on the single solvent dataset. Mean squared error (MSE) and negative log predictive density (NLPD) are averaged across all leave-one-out data splits. We include the shortest path kernel (sp) and the exponential shortest path kernel (esp).
@@ B.1 Reactor details
[…]
@@ B.1.1 Methods
[…]
@@
All the data-points recorded were reported in the dataset, and the only outliers that were removed were those slugs that experienced a step-up in flow-rate while in the reactor, as this has been shown to add bias to the data [schrecker2024comparative].
@@ B.2 Fine-tuning calibration via optimization
$y_{product}=\epsilon_{product}\times\frac{c_{IS}}{c_{0}}\times\text{peak\_ratio}$ (5)
@@
[…]
@@
  • $\text{peak\_ratio}_{ij}$ are the observed HPLC peak area ratios,
@@ B.3 Spange descriptor interpolation
The descriptors from spange2021reappraisal were obtained from the supplementary material on the paper. However, there are a few values missing from some rows, including for the solvents we gathered data for. In order to estimate the missing values, we trained a multi-task Gaussian process model on the whole table, under a Taniamoto kernel, which we then used to predict the missing values that are used for all the main methods in the paper.